Hi, I just took a test that had 10 questions, 8 questions were 2pts each and the other 2 were 7pts, so a total of 30pts. I want to calculate what my grade would be based on a 100pt scale, but I can't remember how. So lets say I got 25 of the 30pts, what would my grade be based on a 100pt scale? Even though I'll only get up to 30pts for this test, whenever I get my grade back I want to be able to tell if I made an A,B,C, and if it would calculate to a 90,80,70... Hope that makes sense! Thanks for the help.
Answer by stanbon(75887)

25/30 = x/100
5/6 = x/100
x = 500/6 = 83.333 or 83 1/3 %
================
Cheers,
Stan H.